# MUSES: The Multi-Sensor Semantic Perception Dataset for Driving under Uncertainty

### v2

@@ 6 Conclusion
We have presented MUSES, the first multi-sensor semantic perception dataset for autonomous driving in adverse conditions. MUSES provides 2500 samples, each consisting of synchronized and calibrated RGB images, MEMS lidar scans, HD event sequences, FMCW radar scans, IMU/GNSS readings, and an image-level corresponding normal-condition image. Our new two-stage panoptic annotation protocol yields high-quality 2D panoptic annotations for each image and class- and instance-level difficulty maps, which enable the novel task of uncertainty-aware panoptic segmentation. Our experiments highlight the benefits of the additional non-camera modalities for dense semantic perception and prove that MUSES is effective for training and provides a challenging evaluation for models under diverse visual conditions. This dataset motivates many possible lines of research: sensor fusion for dense semantic perception tasks, the exploration of strengths and weaknesses of individual sensors, e.g., event cameras, for semantic perception in diverse visual conditions, and model generalization and/or adaptation across conditions or modalities.
@@ Appendix 0.C Calibration
Geometric calibration. The intrinsic parameters for the lidar and radar are vendor-calibrated. For the event camera, we reconstruct frames following [32] and use the same intrinsic calibration procedure as for the frame camera, using a metrology-grade checkerboard and OpenCV [7]. For the extrinsic calibration, we create a consistent transform graph between frame camera, event camera, and lidar: First, we perform pairwise calibrations: Stereo-calibration between the frame and event camera according to [32], mutual information maximization between lidar and frame camera according to [34], and mutual information maximization between lidar and event camera according to [46]. Next, we formulate a triangular pose graph and optimize it for loop closure (Powell’s dog leg method) to get a consistent set of extrinsics.
For radar-lidar calibration we follow [9]: the rotation is estimated via correlative scan matching using a Fourier Mellin transform [12] and the translation is simply measured. The IMU/GNSS is calibrated with u-center [1] and subsequent point cloud consistency optimization.
@@ Appendix 0.D Anonymization
Appendix 0.E Training Details
0.E.1 Experiment Implementation Details
@@
0.E.2 Multimodal Architecture Details
@@
0.E.3 Mask-Classification Baseline for Uncertainty-Aware Panoptic Segmentation
@@
Appendix 0.F Stage 1 vs. Stage 2: Detailed Results
As mentioned in Sec. 5.1 of the main paper, we try to quantify the added difficulty—and implied subsequent quality—of our second labeling stage onto annotations. We train frame-camera-only Mask2Former with a Swin-T backbone on our final panoptic ground truth (equivalent to H2 labels) and evaluate it on stage 1 (H1) and stage 2 (H2) labels. The results by condition presented in Tab. 8 show a 11.1% drop from H2 to H1, indicating substantially more difficult ground truth after the second stage of the annotation. Together with the good generalization results in Sec. 5.3, this indicates that the additional labeled portion of the images (see Fig. 7(a)) is accurately labeled. As these are also more difficult-to-predict areas, these labels guide models effectively during training, leading to improved generalization (see Sec. 5.3).
@@
Appendix 0.G Detailed Class-Level Results
For the models presented in Sec. 5.2, we present detailed class-level PQ results in Tabs. 9, 16, 17, 18, 19, 20 and 21. Adding an event camera to the frame-based camera performs well across the board, with the event camera showing large improvement on small dynamic classes like “motorcycle”, “bicycle” and “person”. Radar seems to be specifically good for larger metallic objects like “bus”, “train” and “car”. Their metal parts make them easier to detect with radar, due to their large radar cross-section. The lidar is generally very helpful in identifying large continuous objects like “building” and “truck”. These larger objects usually have good lidar returns, even in more challenging weather conditions. Further, we can also observe an especially large gap of 7.3% PQ from lidar to the other modalities for “sign”. This is because traffic signs, which have reflective coatings, give off high-intensity readings for lidar.
@@ Appendix 0.G Detailed Class-Level Results
Appendix 0.H Further Dataset Statistics
0.H.1 Things-Classes Statistics
@@
0.H.2 Lidar Point Cloud Statistics
@@ 0.H.3 Difficulty Map Distribution
Appendix 0.I Visualization of MUSES Samples
@@
Appendix 0.J Qualitative Results
We visualize some exemplary panoptic segmentation predictions results of the uni- and quadrimodal Mask2Former with their respective class and instance uncertainty maps in Figs. 14 and 15. The class and instance uncertainty scores are calculated according to Sec. 0.E.3.

### v3

@@ 6 Conclusion
We have presented MUSES, the first multi-sensor semantic perception dataset for autonomous driving in adverse conditions. MUSES provides 2500 samples, each consisting of synchronized and calibrated RGB images, MEMS lidar scans, HD event sequences, FMCW radar scans, IMU/GNSS readings, and an image-level corresponding normal-condition image. Our new two-stage panoptic annotation protocol yields high-quality 2D panoptic annotations for each image and class- and instance-level difficulty maps, which enable the novel task of uncertainty-aware panoptic segmentation. Our experiments highlight the benefits of the additional non-camera modalities for dense semantic perception and prove that MUSES is effective for training and provides a challenging evaluation for models under diverse visual conditions. acrossThis dataset motivates many possible lines of research: sensor fusion for dense semantic perception tasks, the exploration of strengths and weaknesses of individual sensors, e.g., event cameras, for semantic perception in diverse visual conditions, and model generalization and/or adaptation across conditions or modalities.
@@ Appendix 0.C Calibration
Geometric calibration. The intrinsic parameters for the lidar and radar are vendor-calibrated. For the event camera, we reconstruct frames following [33] and use the same intrinsic calibration procedure as for the frame camera, using a metrology-grade checkerboard and OpenCV [7]. For the extrinsic calibration, we create a consistent transform graph between frame camera, event camera, and lidar: First, we perform pairwise calibrations: Stereo-calibration between the frame and event camera according to [33], mutual information maximization between lidar and frame camera according to [35], and mutual information maximization between lidar and event camera according to [48]. Next, we formulate a triangular pose graph and optimize it for loop closure (Powell’s dog leg method) to get a consistent set of extrinsics.Author For radar-lidar calibration we follow [9]: the rotation is estimated via correlative scan matching using a Fourier Mellin transform [12] and the translation is simply measured. The IMU/GNSS is calibrated with u-center [1] and subsequent point cloud consistency optimization.
@@ Appendix 0.D Anonymization
Appendix 0.E Uncertainty Type
Assuming human annotation as the gold standard—given our intensive quality control—any human uncertainty is attributed to aleatoric uncertainty (data-inherent and irreducible). “Difficult” pixels are those explained only by additional data available at annotation stage 2 (e.g., videos, lidar). Hence, a model that only has access to stage-1 (camera) data should not be confident in predicting the labels of “difficult” pixels, as their semantics cannot be explained by this data alone. UPQ allows the model to acknowledge uncertainty alongside making a correct prediction, thus encouraging an uncertainty-aware prediction. This rationale only holds for UPQ evaluation of camera-based models, and not multimodal ones.
Appendix 0.F Training Details
0.F.1 Experiment Implementation Details
@@
0.F.2 Multimodal Architecture Details
@@
0.F.3 Mask-Classification Baseline for Uncertainty-Aware Panoptic Segmentation
@@
Appendix 0.G Stage 1 vs. Stage 2: Detailed Results
As mentioned in Sec. 5.1 of the main paper, we try to quantify the added difficulty—and implied subsequent quality—of our second labeling stage onto annotations. We train frame-camera-only Mask2Former with a Swin-T backbone on our final panoptic ground truth (equivalent to H2 labels) and evaluate it on stage 1 (H1) and stage 2 (H2) labels. The results by condition presented in Tab. 8 show a 11.1% drop from H2 to H1, indicating substantially more difficult ground truth after the second stage of the annotation.
To further investigate the quality of the additional annotations, we train a semantic segmentation Mask2Former (Swin-L) exclusively on H1 annotations. The performance of this model is shown in Tab. 9. The model trained on H1 annotations shows a significant performance drop on ACDC and MUSES datasets. Specifically, there is a drop of 3.3 mIoU on MUSES and 2.0 mIoU on ACDC compared to the model trained on H2 annotations. On the other hand, the scores on Cityscapes are similar for both H1 and H2-trained models. This result is expected because the additional H2 labels primarily address indistinct areas in adverse scenes, which are not present in Cityscapes. Cityscapes labels do not include adverse conditions or auxiliary data, making them more aligned with the H1 labels. As a result, the performance of the H1 and H2 models on Cityscapes is similar. In contrast, the ACDC dataset, which includes some auxiliary data, benefits significantly from the higher-quality H2 labels.
These findings collectively underscore the high quality of the additional H2 labels, as they enhance model performance in challenging conditions where auxiliary data is crucial. Together with the good generalization results in Sec. 5.3, these results indicate that the additional labeled portion of the images (see Fig. 7 left) is accurately labeled. As these are also more difficult-to-predict areas, these labels guide models effectively during training, leading to improved generalization (see Sec. 5.3).
@@
Table 9: Mask2Former [15] (Swin-L [29]) performance trained on H2 versus H1 annotations.
mIoU $\uparrow$ Cityscapes ACDC MUSES
Train on MUSES–H2 73.1 72.0 77.1
Train on MUSES–H1 73.3 70.0 73.8
Appendix 0.H Detailed Class-Level Results
For the models presented in Sec. 5.2, we present detailed class-level PQ results in Tabs. 10, 17, 18, 19, 20, 21 and 22. Adding an event camera to the frame-based camera performs well across the board, with the event camera showing large improvement on small dynamic classes like “motorcycle”, “bicycle” and “person”. Radar seems to be specifically good for larger metallic objects like “bus”, “train” and “car”. Their metal parts make them easier to detect with radar, due to their large radar cross-section. The lidar is generally very helpful in identifying large continuous objects like “building” and “truck”. These larger objects usually have good lidar returns, even in more challenging weather conditions. Further, we can also observe an especially large gap of 7.3% PQ from lidar to the other modalities for “sign”. This is because traffic signs, which have reflective coatings, give off high-intensity readings for lidar.
@@ Appendix 0.H Detailed Class-Level Results
Appendix 0.I Further Dataset Statistics
0.I.1 Things-Classes Statistics
@@
0.I.2 Lidar Point Cloud Statistics
@@ 0.I.3 Difficulty Map Distribution
Appendix 0.J European Domain Bias
The MUSES dataset consists exclusively of driving scenes from Switzerland, which may introduce a geographical bias towards Western European environments. This limitation is inherent in the dataset’s design and location. While this regional focus ensures consistency and depth within a specific context, it may affect the generalizability of the results to other regions with different driving conditions, infrastructure, and weather patterns.
Appendix 0.K Visualization of MUSES Samples
@@
Appendix 0.L Qualitative Results
We visualize some exemplary panoptic segmentation predictions results of the uni- and quadrimodal Mask2Former with their respective class and instance uncertainty maps in Figs. 14 and 15. The class and instance uncertainty scores are calculated according to Sec. 0.F.3.